Meeting Time and Place
The meeting will be on Wednesday July 24, 2013 at 11:15am EDT. For those people at Jefferson Lab, the meeting will be in room F326.
To connect from the outside:
Call ESNET Number 8542553 (this is the preferred connection method).
2.) Polycom RealPresence App on iOS and Android:
3.) Phone: (should not be needed)
+1-866-740-1260 : US and Canada +1-303-248-0285 : International then use participant code: 3421244# (the # is needed when using the phone) or www.readytalk.com then type access code 3421244 into "join a meeting" (you need java plugin)
- Solenoid Status
MOU and Contract Status
- MOU Status
- UConn (Diamond)
Reports from Technical Groups
- Tracking July 17, 2013
- Engineering Activities
- Installation July 19, 2013 Installation
- Technical Design Report
- Trigger July 17, 2013
- DAQ/Monitoring/Controls July 17, 2013
- Offline June 26, 2013
- Photon Beam and Tagger July 22, 2013
- Physics June 17, 2013
Present: MIT, FSU, CMU, Regina, UConn, IU, JLab
Solenoid: Trying to cool the solenoid but difficult. Refrigerator is not working well. The short is gone. Power supply has been improved. Everything is ready to go, but waiting for cooling. Two dewars are being supplied every day for cooling. There may be limitations on the maximum current. Nothing above 1350A is possible and it may be lower.
Contracts: FIU is done. No one knows the status of the UConn MOU for diamond radiator. It is in the hands of Jefferson Lab. Elton will check on it.
FCAL: 1/4 is cabled. Work continues on the environmental monitoring system.
BCAL: Building the 49th module now. Modules will be moved to the hall for assembly once the work with the magnet is done.
Calorimeter monitoring: Sheets of UVT Plexi for the FCAL monitoring system are hard to find. Sheets that are 1/2" that are larger than 48" in both dimensions are very difficult to get. We may need to fabricate new clamps to accommodate smaller sheets.
CDC: Naomi is having some success with CDC readout using the VXS crate. Cooling rings are in place and Gary is working on support structure for HV connection.
FDC: Beni completed test of fully cabled package. Noise levels are low. One bad flex board connection needed to be fixed. The full package has no dead channels. Two packages have been mounted together already and eventually all four will be connected and moved to EEL by the end of the month.
TOF: Production continues at a rate of about 9 modules a week. Issue with PMTs in gluing. 10% have a dull spot due to surface imperfections on the light guide that resulted from bad polishing. Some tubes have been removed and light guides are now being polished again to flatten the surface. There is a slight delay, but still plan to be finished ahead of schedule by the end of August or early September.
Start Counter: no report
PID Upgrade: Two page document is available on portal that summarizes the gDIRC concept. The goal is to plan a trip to SLAC to inspect the bars and understand what is needed to transport and refurbish them.
Trigger: Trigger tests continue. In August we expect CTP modules and after they are checked they will be used for the BCAL cosmic ray tests.
L3 Trigger: Hoping for a preliminary L3 algorithm for online data challenge.
DAQ/Monitoring: Online data challenge is planned for the last week of August that will include data transfer, monitoring, event rejection and the whole online data stream.
Controls: CDC high voltage control is finished. FDC is partially finished. Next tasks will be the low voltage systems.
We need to develop the specifications for the RF signal that the accelerator group will provide to Hall D. The beamline group will address this in their next meeting.
Offline: Plan to discuss the outcome of the tutorial this afternoon. Mark is developing a list of tasks for the online data challenge.
Beamline and Tagger: Microscope fabrication continues smoothly planning an on schedule delivery. Electronics has been received at Jefferson Lab. Preliminary estimate from ASU on triplet polarimetry detector is about $50K. We are figuring out how to fund this. Assembly of low granularity counters is going well at North Carolina. Studies are underway to understand how finely the tagger mapping needs to be done in the fringe field of the tagger magnet. The microscope needs to be moved to accommodate the pion polarizability experiment -- need to understand how to reroute cables.
Physics: Meeting on Monday to compile a list of projects and reactions to study in order to study the physics and offline reconstruction performance. Expect enhanced capabilities after the analysis workshop.